If the potential difference is applied across a resistor of length l and diameter d .if the diameter will halfed then the V will be


1
Expert's answer
2021-06-28T17:06:03-0400

Gives

Patentialdifferrnce=V

resistor=R

Length=l

Diameter (d1)=d

We know that


"R_1=\frac{\rho l}{A_1}=\frac{\rho l_1}{\pi d_1^2}"

"V=IR_1"

"V_1=I_1\frac{\rho l_1}{\pi d_1^2}"

When

"d_2=\frac{d}{2}
,l_2=2l"

"R_2=\frac{\rho l_2
}{A_2}"

"R_2=8\frac{\rho l_1}{\pi d_1^2}"

"V_2=IR_2" ";V_2=8I \frac{\rho l}{\pi d_1^2}"

"V_2=8V_1"